[Brahms-l] Shift report 20040115 16:00-24:00

From: Apache <apache@bnl.gov>
Date: Thu Jan 15 2004 - 23:47:27 EST
------------------------------------------------------------
Date: Thu Jan 15 23:47:27 2004, Shift: 16:00-24:00
Supervisor: ZIS, with: 
Log book#15 pages 33 - 33

Shift Summary:
------------------------------------------------------------
1. Store from 20:30 until 23:00.
- FS and BS has been rotated at 12 degrees
- the magnetic field has been decreased to: 
<li> B430
<p>
2. Runs:
<li> 9701 ( 312 K);
<li> 9702 (  65 K);
<li> 9704 (  82 K);
<p>
3. Detector problems:
<li> The refrigerator for the alchol bubbler has failed. 
<li> The zdc trigger was adjusted to reduce the vertex selected for triggers
<li> Hiro and Bob went in IR to fix Tile HV (4032 mainframe 5), TOFW 1440 (#1) and T5 low voltage

<p>
<a href="http://pii3.brahms.bnl.gov/cgi-bin/elog.pl?nb=2004&action=view&page=43&scroll=false"> more info see E-log</a>

------------------------------------------------------------

_______________________________________________
Brahms-l mailing list
Brahms-l@lists.bnl.gov
http://lists.bnl.gov/mailman/listinfo/brahms-l
Received on Thu Jan 15 23:47:47 2004

This archive was generated by hypermail 2.1.8 : Thu Jan 15 2004 - 23:48:08 EST